# Look Before You Leap: Problem Elaboration Prompting Improves Mathematical Reasoning in Large Language Models

### v1

@@ Appendix A Prompt Details
The other prompts used in our experiments.
@@
All prompts are placed before the given problem. Then we extracted the results from outputs based on regular expressions. However, the outputs may still be diverse and not compliant with the formatting. Therefore, we manually compared and extracted all the non-numeric and non-option items.
Appendix B Example Input and Output Pairs
Q: There were 7 roses in the vase. Jessica cut some more roses from her flower garden which had a total of 59 roses. There are now 20 roses in the vase. How many roses did she cut?
Standard CoT (Solve the question step by step.):
A: To solve the problem, we need to find out how many roses Jessica cut from her garden.
We know that there were originally 7 roses in the vase, and now there are 20. So Jessica must have added 20 - 7 = 13 roses to the vase.
We also know that Jessica cut these roses from her flower garden, which had a total of 59 roses. So the number of roses she cut is: 59 - 13 = 46
Therefore, Jessica cut 46 roses from her garden. (wrong)
PEP + CoT (Decompose the given question into smaller segments, elucidating each segment as you rephrase it. Then, solve the question step by step.):
Segment 1: How many roses were in the vase initially?
- There were 7 roses in the vase.
Segment 2: How many roses did Jessica cut from her flower garden?
- Jessica cut some more roses from her flower garden.
Segment 3: How many roses were in Jessica’s flower garden initially?
- Jessica’s flower garden had a total of 59 roses.
Segment 4: How many roses are in the vase now?
There are now 20 roses in the vase.
@@ Appendix B Example Input and Output Pairs
To find out how many roses Jessica cut from her flower garden, we need to subtract the initial number of roses in the vase from the current number of roses in the vase.
Initial number of roses in the vase = 7
Current number of roses in the vase = 20
Number of roses Jessica added to the vase = Current number of roses in the vase - Initial number of roses in the vase= 20 - 7 = 13
Therefore, Jessica cut 13 roses from her flower garden.
Table 12: The complete example of Fig. 1 from SVAMP, generated by turbo with greedy decoding. The sentence (cyan) is complex and confusing as the misleading condition is mixed with the aiming condition. The standard CoT does not correctly parse the problem (red), while PEP  (blue) benefits from the decomposition and elucidation to clarify the problem.
Question 1:
@@
It is approximately 1955 kilometers from San Diego, California to New York City, New York. If Bernice drove 325 kilometers for 4 days, how many kilometers will she still need to drive?
CoT
@@
To find the total distance Bernice drove, we can multiply the distance she drove each day (325 kilometers) by the number of days she drove (4 days): 325 km/day x 4 days = 1300 kilometers. Therefore, Bernice drove 1300 kilometers. To find how many kilometers she still needs to drive, we can subtract the distance she drove from the total distance between San Diego and New York City: 1955 km - 1300 km = 655 kilometers. Therefore, Bernice still needs to drive 655 kilometers. (Correct)
@@
PEP + CoT:
Segment 1: What is the distance between San Diego, California and New York City, New York? The distance between San Diego, California and New York City, New York is approximately 1955 kilometers. Segment 2: How many kilometers did Bernice drive? Bernice drove 325 kilometers. Segment 3: How many days did Bernice drive for? Bernice drove for 4 days. Segment 4: How many kilometers will Bernice still need to drive? To find out how many kilometers Bernice will still need to drive, we need to subtract the distance she has already driven from the total distance between San Diego and New York. 1955 km - 325 km = 1630 km. Therefore, Bernice will still need to drive 1630 kilometers.
Question 2:
Sandy wants to lose as much weight as Joey does but needs 4 weeks to lose the same amount of weight that Joey loses in a single week. If Joey loses 8 pounds in 4 weeks, how many weeks will it take Sandy to lose the same amount of weight?
CoT
First, we need to find out how much weight Joey loses in a single week. We can do this by dividing the total weight he loses in 4 weeks by 4: 8 pounds ÷ 4 weeks = 2 pounds per week. So Joey loses 2 pounds per week. Next, we need to figure out how much weight Sandy wants to lose. We know she wants to lose the same amount as Joey, which is 8 pounds. Now we can use the information given in the problem to find out how many weeks it will take Sandy to lose 8 pounds. We know that Sandy needs 4 weeks to lose the same amount of weight that Joey loses in a single week. So we can set up a proportion: 4 weeks to lose 2 pounds = x weeks to lose 8 pounds. To solve for x, we can cross-multiply: 4 weeks * 8 pounds = 2 pounds * x weeks; 32 pounds = 2x; Divide both sides by 2: 16 weeks = x. So it will take Sandy 16 weeks to lose the same amount of weight that Joey loses in 4 weeks. (Correct)
@@
PEP + CoT:
Segment 1: Sandy wants to lose as much weight as Joey does. Elucidation: Sandy has a weight loss goal that is equivalent to Joey’s weight loss goal. Segment 2: Sandy needs 4 weeks to lose the same amount of weight that Joey loses in a single week. Elucidation: Sandy’s weight loss rate is slower than Joey’s weight loss rate. It takes Sandy 4 weeks to lose the same amount of weight that Joey loses in just one week. Segment 3: Joey loses 8 pounds in 4 weeks. Elucidation: Joey’s weight loss goal is to lose 8 pounds, and he achieves this goal in 4 weeks. Segment 4: How many weeks will it take Sandy to lose the same amount of weight? Elucidation: The question is asking for the number of weeks it will take Sandy to achieve the same weight loss goal as Joey.
Solution: Since Joey loses 8 pounds in 4 weeks, his weight loss rate is 8/4 = 2 pounds per week. Since Sandy needs 4 weeks to lose the same amount of weight that Joey loses in a single week, her weight loss rate is 1/4 = 0.25 pounds per week. To lose 8 pounds like Joey, Sandy will need to lose at the same rate of 2 pounds per week. Therefore, it will take Sandy 8/2 = 4 weeks to lose the same amount of weight as Joey.
Question 3:
Patty’s Plumbing charges \$40 to visit a house to make a repair, plus \$35 per hour, or part thereof, for labor, plus parts. One job took 2.25 hours and used \$60 in parts. How much did Patty charge?
CoT
1. Calculate the labor cost: The job took 2.25 hours, so Patty charged for 3 hours (since she charges for any part of an hour). The labor cost is \$35 per hour, so the labor cost for 3 hours is 3 x \$35 = \$105. 2. Calculate the total cost: The visit fee is \$40. The labor cost is \$105. The parts cost is \$60. The total cost is \$40 + \$105 + \$60 = \$205. Therefore, Patty charged \$205 for the job. (Correct)
@@
PEP + CoT:
Segment 1: What is the base fee for Patty’s Plumbing to visit a house for a repair? Patty’s Plumbing charges \$40 to visit a house for a repair. Segment 2: How much does Patty’s Plumbing charge for labor? Patty’s Plumbing charges \$35 per hour, or part thereof, for labor. Segment 3: How much does Patty’s Plumbing charge for parts? The job used \$60 in parts. Segment 4: How long did the job take? The job took 2.25 hours.
Solution: To calculate the total cost of the job, we need to add the base fee, labor cost, and parts cost. Base fee: \$40 Labor cost: \$35 per hour x 2.25 hours = \$78.75 Parts cost: \$60. Total cost: \$40 + \$78.75 + \$60 = \$178.75. Therefore, Patty charged \$178.75 for the job.
@@
The complete results and comparison of Fig. 1 is shown in Table 12. The given sentence in the problem is complex and confusing, as it mixed the condition "Jessica cut some more roses" which is associated with the final question and the irrelevant sentence "which had a total of 59 roses". CoT is easily misled, while PEP correctly decomposes these conditions into two segments and elucidate the behind meanings, where the downstream reasoning process gets more clear and coherent.

### v2

@@ Appendix A Experiment Details
Model Usage.
We used turbo-0301, an early version of ChatGPT, for instruction selection and type analysis in PEP. Given the possibility that OpenAI may restrict access to early models, we tested PEP on four open-source models, text-davinci-003 and the recently released turbo-0125, as well to validate the generalization of the instructions. We utilized both Llamma models and Mistral-7B with bfloat16. Due to cost constraints, we loaded Mistral-8x7B in 4 bits.
Appendix B All used instructions and exemplars
B.1 Zero-shot prompts
To ensure fairness and clarity of semantics when combining multiple instructions, we use “Let’s solve the question step by step” as the zero-shot CoT instead of “Let’s think step by step.”. We compared these two instructions in Table 9. The used prompts are list as follow:
@@ B.1 Zero-shot prompts
[…]
IRR_Inst = ”Feel free to ignore irrelevant information given in the questions.”
B.2 Extract the answers
In order to extract results better, we add a standardized output instruction after each prompt during generation, namely
(1) for free-answered questions, we use
[…]
(2) for questions with options, we use
[…]
Finally, we use a one-shot exemplars to extract the answers from generations by turbo-0125. For those unrecognized solutions, we extract the answers manually.
Extract_Template = “‘Given the textual solution or code execution solution, output the numeric answer that can be converted into float value of the problem. If the solution does not yield a result, output ”unsolved”. Only output the numeric value or ”unsolved”.
### Example:
Solution: The total amount of money Janet makes from selling eggs at the farmers’ market per day is 21 - 4 = 17 eggs x \$2 = \$34.
Therefore, the final answer is: Janet makes \$34 per day at the farmers’ market.
Answer: \$34
Result: 34
Original Problem: {qst}
Solution: {sol}
Result:“‘
B.3 Few-shot exemplars
In this section, we enumerate all the exemplars used. They are directly adopted for CoT and L2M from Zhou et al. (2022a), which are specifically designed for GSM8k. For PEP, PaS, and Self-ask, we re-generate the exemplars using GPT-4 for proper modifications.
In practice, the one-shot setting uniformly employs the first exemplar, while the four-shot utilizes all examples. When integrating PEP with other problem-related methods, we simply append the corresponding elaboration part of PEP to the original question of other examples, before any other generations. When employing IRR-Inst, we position it at the beginning.
The overall template structure is as follows, using the one-shot L2M+PEP+IRR_inst as an example for illustration:
[…]
Problem Elaboration:
Segment 1: Elsa has 5 apples.
Segment 2: Anna has 2 more apples than Elsa.
Segment 3: How many apples do they have together?
Rephrased question: If Elsa has 5 apples and Anna has 2 more apples than Elsa, how many apples do they have together?
Answer: Let’s break down this rephrased problem: 1. How many apples does Anna have? 2. How many apples do Elsa and Anna have together?
1. Anna has 2 more apples than Elsa. So Anna has 2 + 5 = 7 apples.
2. Elsa and Anna have 5 + 7 = 12 apples together.
Question: {qst}
Problem Elaboration:“‘
B.3.1 exemplars for CoT
Question: Elsa has 5 apples. Anna has 2 more apples than Elsa. How many apples do they have together?
Answer: Anna has 2 more apples than Elsa, so Anna has 2 + 5 = 7 apples. Elsa and Anna have 5 + 7 = 12 apples together. The answer is 12.
Question: Four years ago, Kody was only half as old as Mohamed. If Mohamed is currently twice 30 years old, how old is Kody?
Answer: We were told that Mohamed is currently twice 30 years old, so he is currently 30 * 2 = 60 years old. That means that four years ago he must have been 60 - 4 = 56 years old. Four years ago, Kody was half as old as Mohamed, so Kody must have been 56 / 2 = 28 years old then. Since Kody was 28 years old four years ago, she must now be 28 + 4 = 32 years old. The answer is 32.
Question: Carla bought 2 bags of mini peanut butter cups on clearance. Each bag was \$6.00 but was 75% off. How much did she spend on 2 bags of candy?
Answer: Each bag was \$6.00 but was 75% off. So each bag cost \$6.00 * (1 - 0.75) = \$6.00 * 0.25 = \$1.50. Carla bought 2 bags. So she spent \$1.50 * 2 = \$3.00. The answer is 3.
Question: If Pam is currently twice as young as Rena is, and in 10 years Rena will be 5 years older than her, how old is Pam now?
Answer: Since Rena will be 5 years older than Pam in 10 years, she must be 5 years older than Pam now as well. If Pam is currently twice as young as Rena, that means that Rena is currently twice as old as Pam is. So if P stands for Pam’s age now and R stands for Rena’s age now, then we know that R = 2 * P And since Rena is 5 years older than Pam now, we know that R = P + 5. By substitution, we have P + 5 = 2 * P, which means that P = 5. The answer is 5.
FOUR_SHOT_CoT = “‘{exemplars}
Question: {qst}
Answer:
B.3.2 exemplars for PEP
Question: Elsa has 5 apples. Anna has 2 more apples than Elsa. How many apples do they have together?
Problem Elaboration:
Segment 1: Elsa has 5 apples. This segment tells us the number of apples Elsa has.
Segment 2: Anna has 2 more apples than Elsa. This segment tells us that Anna has more apples than Elsa, specifically 2 more.
Segment 3: How many apples do they have together? This segment is asking us to find the total number of apples both Elsa and Anna have combined.
@@ B.3 Few-shot exemplars
Step 1: Determine the number of apples Elsa has. Elsa has 5 apples.
Step 2: Determine the number of apples Anna has. Anna has 2 more apples than Elsa, so she has 5 + 2 = 7 apples.
Step 3: Determine the total number of apples they have together. Together, Elsa and Anna have 5 + 7 = 12 apples. The answer is 12.
Question: Four years ago, Kody was only half as old as Mohamed. If Mohamed is currently twice 30 years old, how old is Kody?
Problem Elaboration:
Segment 1: Four years ago, Kody was only half as old as Mohamed. This means that the age difference between Kody and Mohamed is constant and it is the same four years ago as it is now.
Segment 2: If Mohamed is currently twice 30 years old. This means that Mohamed’s current age is 60 years old.
Solution:
Step 1: Determine Mohamed’s age four years ago. If Mohamed is currently 60 years old, then four years ago he was 60 - 4 = 56 years old.
Step 2: Determine Kody’s age four years ago. Since Kody was half as old as Mohamed four years ago, then Kody was 56 / 2 = 28 years old four years ago.
Step 3: Determine Kody’s current age. If Kody was 28 years old four years ago, then Kody is currently 28 + 4 = 32 years old. The answer is 32.
Question: Carla bought 2 bags of mini peanut butter cups on clearance. Each bag was \$6.00 but was 75% off. How much did she spend on 2 bags of candy?
Problem Elaboration:
Segment 1: Identify the original price of the bags of candy. The original price of each bag of candy is \$6.00.
Segment 2: Determine the discount on each bag. The bags are 75% off.
Segment 3: Calculate the discounted price of each bag. To find the discounted price, we need to calculate 75% of \$6.00.
Segment 4: Determine the total cost for 2 bags. Once we have the discounted price of one bag, we multiply it by 2 to find the total cost for 2 bags.
Solution: Step 1: The original price of each bag is \$6.00. Step 2: The discount on each bag is 75%. Step 3: To calculate 75% of \$6.00, we multiply 6 by 0.75, which equals \$4.50. This means that \$4.50 is the amount of the discount. Step 4: To find the discounted price of each bag, we subtract the discount from the original price. So, \$6.00 - \$4.50 = \$1.50. Each bag costs \$1.50 after the discount. Step 5: To find the total cost for 2 bags, we multiply the discounted price by 2. So, \$1.50 * 2 = \$3.00. The answer is 3.
Question: If Pam is currently twice as young as Rena is, and in 10 years Rena will be 5 years older than her, how old is Pam now?
Problem Elaboration:
Segment 1: Pam is currently twice as young as Rena is. This means that Pam’s current age is half of Rena’s current age.
Segment 2: In 10 years, Rena will be 5 years older than Pam. This means that if we add 10 years to both Pam’s and Rena’s current ages, the difference between their ages will be 5 years.
Solution:
Step 1: Let’s denote Rena’s current age as R and Pam’s current age as P. From the first segment, we know that P = R/2.
Step 2: From the second segment, we know that R + 10 = P + 10 + 5. We can simplify this to R = P + 5.
Step 3: Now we can substitute P from the first equation into the second equation. So, R = R/2 + 5.
Step 4: To solve for R, we multiply both sides of the equation by 2 to get rid of the fraction. This gives us 2R = R + 10.
Step 5: Subtract R from both sides to get R = 10. So, Rena is currently 10 years old.
Step 6: Substitute R = 10 into the first equation to find P. This gives us P = 10/2 = 5. So, Pam is currently 5 years old. The answer is 5.
FOUR_SHOT_PEP = “‘{exemplars}
Question: {qst}
Problem Elaboration: “‘
B.3.3 exemplars for L2M
Question: Elsa has 5 apples. Anna has 2 more apples than Elsa. How many apples do they have together?
Answer: Let’s break down this problem: 1. How many apples does Anna have? 2. How many apples do Elsa and Anna have together?
1. Anna has 2 more apples than Elsa. So Anna has 2 + 5 = 7 apples.
2. Elsa and Anna have 5 + 7 = 12 apples together. The answer is 12.
Question: Four years ago, Kody was only half as old as Mohamed. If Mohamed is currently twice 30 years old, how old is Kody?
Answer: Let’s break down this problem: 1. How old was Mohamed four years ago? 2. How old is Kody?
1. We were told that Mohamed is currently twice 30 years old, so he is currently 30 * 2 = 60 years old. That means that four years ago he must have been 60 - 4 = 56 years old.
2. Four years ago, Kody was half as old as Mohamed, so Kody must have been 56 / 2 = 28 years old then. Since Kody was 28 years old four years ago, she must now be 28 + 4 = 32 years old. The answer is 32.
Question: Carla bought 2 bags of mini peanut butter cups on clearance. Each bag was \$6.00 but was 75% off. How much did she spend on 2 bags of candy?
Answer: Let’s break down this problem: 1. How much did she spend on 2 bags of candy?
1. Each bag was \$6.00 but was 75% off. So each bag cost \$6.00 * (1 - 0.75) = \$6.00 * 0.25 = \$1.50. Carla bought 2 bags. So she spent \$1.50 * 2 = \$3.00. The answer is 3.
Question: If Pam is currently twice as young as Rena is, and in 10 years Rena will be 5 years older than her, how old is Pam now?
Answer: Let’s break down this problem: 1. How much older is Rena than Pam currently? 2. How old is Pam now?
1. Since Rena will be 5 years older than Pam in 10 years, she must be 5 years older than Pam now as well.
2. If Pam is currently twice as young as Rena, that means that Rena is currently twice as old as Pam is. So if P stands for Pam’s age now and R stands for Rena’s age now, then we know that R =2 * P And since Rena is 5 years older than Pam now, we know that R = P + 5. By substitution, we have P + 5 = 2 * P, which means that P = 5. The answer is 5.
FOUR_SHOT_L2M = “‘{exemplars}
Question: {qst}
Answer: Let’s break down this problem:“‘
B.3.4 exemplars for Self-ask
Question: Elsa has 5 apples. Anna has 2 more apples than Elsa. How many apples do they have together?
Are follow up questions needed here: Yes.
Follow up: How many apples does Anna have?
Intermediate answer: Anna has 5 + 2 = 7 apples.
Follow up: How many apples do Elsa and Anna have together?
Intermediate answer: Elsa and Anna have 5 + 7 = 12 apples together. The answer is 12.
Question: Four years ago, Kody was only half as old as Mohamed. If Mohamed is currently twice 30 years old, how old is Kody?
Are follow up questions needed here: Yes.
Follow up: How old is Mohamed currently?
Intermediate answer: Mohamed is 30 * 2 = 60 years old. Follow up: How old was Kody four years ago?
Intermediate answer: Kody was (60 - 4) / 2 = 28 years old four years ago.
So the final answer is: Kody is 28 + 4 = 32 years old. The answer is 32.
Question: Carla bought 2 bags of mini peanut butter cups on clearance. Each bag was \$6.00 but was 75% off. How much did she spend on 2 bags of candy?
Are follow up questions needed here: Yes.
Follow up: How much was the discount for each bag?
Intermediate answer: The discount for each bag is \$6.00 * 75% = \$4.50.
Follow up: How much did Carla pay for each bag after the discount?
Intermediate answer: Carla paid \$6.00 - \$4.50 = \$1.50 for each bag.
So the final answer is: Carla spent \$1.50 * 2 = \$3.00 on 2 bags of candy. The answer is 3.00.
Question: If Pam is currently twice as young as Rena is, and in 10 years Rena will be 5 years older than her, how old is Pam now?
Are follow up questions needed here: Yes.
Follow up: What about Rena and Pam’s current ages?
Intermediate answer: It tells us that Rena’s age is twice Pam’s age. So if P stands for Pam’s age now and R for Rena’s age now, then R = 2 * P. And since Rena is 5 years older than Pam now, we have R = P + 5.
Follow up: What is Pam’s age now?
Final answer: By substituting P + 5 in place of R in equation R = 2 * P, we get P + 5 = 2 * P, which simplifies to P = 5. So, Pam is 5 years old. The answer is 5.
FOUR_SHOT_SK =“‘{exemplars}
Question: {qst}
Are follow up questions needed here: Yes.
Follow up: “‘
It’s worth noting that the template recommended in self-ask (Press et al., 2022) actually ends with “Are follow up questions needed here:”, but we found it always generates a “Yes” or “No” and then stops, especially in one-shot setting. Therefore, we made above adjustments.
B.3.5 exemplars of PaS
Question: Elsa has 5 apples. Anna has 2 more apples than Elsa. How many apples do they have together?
Plan:
Step 1: Find out how many apples Anna has.
Step 2: Add the number of apples Elsa has to the number of apples Anna has to find the total number of apples they have together.
Solution:
Step 1: Anna has 5 + 2 = 7 apples.
Step 2: Together, Elsa and Anna have 5 (Elsa’s apples) + 7 (Anna’s apples) = 12 apples. So, Elsa and Anna have 12 apples together. The answer is 12.
Question: Four years ago, Kody was only half as old as Mohamed. If Mohamed is currently twice 30 years old, how old is Kody?
Plan:
Step 1: Find out how old Mohamed is currently.
Step 2: Find out how old Mohamed was four years ago.
Step 3: Since Kody was half as old as Mohamed four years ago, find out Kody’s age four years ago.
Step 4: Add four years to Kody’s age to find out his current age.
Solution:
Step 1: Mohamed is currently 2 * 30 = 60 years old.
Step 2: Four years ago, Mohamed was 60 - 4 = 56 years old.
Step 3: Four years ago, Kody was 56 / 2 = 28 years old. Step 4: Currently, Kody is 28 + 4 = 32 years old. So, Kody is 32 years old. The answer is 32.
Question: Carla bought 2 bags of mini peanut butter cups on clearance. Each bag was \$6.00 but was 75% off. How much did she spend on 2 bags of candy?
Plan:
Step 1: Find out how much discount Carla got on each bag.
Step 2: Subtract the discount from the original price to find out the price Carla paid for each bag. Step 3: Multiply the price Carla paid for each bag by the number of bags she bought to find out how much she spent in total.
Solution:
Step 1: The discount on each bag is 75/100 * \$6.00 = \$4.50.
Step 2: The price Carla paid for each bag is \$6.00 - \$4.50 = \$1.50.
Step 3: Carla spent \$1.50 * 2 = \$3.00 on 2 bags of candy. So, Carla spent \$3.00 on 2 bags of candy. The answer is 3.00.
Question: If Pam is currently twice as young as Rena is, and in 10 years Rena will be 5 years older than her, how old is Pam now?
Plan:
Step 1: Set up an equation based on the information that Rena’s age is twice Pam’s age.
Step 2: Set up another equation based on the information that Rena is 5 years older than Pam.
Step 3: Substitute the second equation into the first to solve for Pam’s age.
Solution:
Step 1: If P stands for Pam’s age now and R for Rena’s age now, then R = 2 * P.
Step 2: And since Rena is 5 years older than Pam now, we have R = P + 5.
Step 3: By substituting P + 5 in place of R in equation R = 2 * P, we get P + 5 = 2 * P, which simplifies to P = 5. So, Pam is 5 years old. The answer is 5.
FOUR_SHOT_PaS = “‘{exemplars}
Question: {qst}
Plan:“‘
B.4 Implementation of PoT
We implement PoT using the following template. For PEP, we insert null into {ela}. For the PoT+PEP, we first use zero-shot PEP to generate the content for the Elaboration part, and then insert it into {ela}.
PoT_zeroshot_temp = “‘# Question: {qst}
{ela}
# Answer the question by implementing a solution() function.
# Generate the code only.
# Let’s write a Python program step by step, and then return the answer
# Firstly, we need write the solution() starting with defining variable:“‘
Appendix C Example Input and Output Pairs
@@
Question 1: It is approximately 1955 kilometers from San Diego, California to New York City, New York. If Bernice drove 325 kilometers for 4 days, how many kilometers will she still need to drive?
@@
CoT To find the total distance Bernice drove, we can multiply the distance she drove each day (325 kilometers) by the number of days she drove (4 days): 325 km/day x 4 days = 1300 kilometers. Therefore, Bernice drove 1300 kilometers. To find how many kilometers she still needs to drive, we can subtract the distance she drove from the total distance between San Diego and New York City: 1955 km - 1300 km = 655 kilometers. Therefore, Bernice still needs to drive 655 kilometers. (Correct)
@@
PEP + CoT: Segment 1: What is the distance between San Diego, California and New York City, New York? The distance between San Diego, California and New York City, New York is approximately 1955 kilometers. Segment 2: How many kilometers did Bernice drive? Bernice drove 325 kilometers. Segment 3: How many days did Bernice drive for? Bernice drove for 4 days. Segment 4: How many kilometers will Bernice still need to drive? To find out how many kilometers Bernice will still need to drive, we need to subtract the distance she has already driven from the total distance between San Diego and New York. 1955 km - 325 km = 1630 km. Therefore, Bernice will still need to drive 1630 kilometers.
Question 2: Sandy wants to lose as much weight as Joey does but needs 4 weeks to lose the same amount of weight that Joey loses in a single week. If Joey loses 8 pounds in 4 weeks, how many weeks will it take Sandy to lose the same amount of weight?
CoT First, we need to find out how much weight Joey loses in a single week. We can do this by dividing the total weight he loses in 4 weeks by 4: 8 pounds ÷ 4 weeks = 2 pounds per week. So Joey loses 2 pounds per week. Next, we need to figure out how much weight Sandy wants to lose. We know she wants to lose the same amount as Joey, which is 8 pounds. Now we can use the information given in the problem to find out how many weeks it will take Sandy to lose 8 pounds. We know that Sandy needs 4 weeks to lose the same amount of weight that Joey loses in a single week. So we can set up a proportion: 4 weeks to lose 2 pounds = x weeks to lose 8 pounds. To solve for x, we can cross-multiply: 4 weeks * 8 pounds = 2 pounds * x weeks; 32 pounds = 2x; Divide both sides by 2: 16 weeks = x. So it will take Sandy 16 weeks to lose the same amount of weight that Joey loses in 4 weeks. (Correct)
@@
PEP + CoT: Segment 1: Sandy wants to lose as much weight as Joey does. Elucidation: Sandy has a weight loss goal that is equivalent to Joey’s weight loss goal. Segment 2: Sandy needs 4 weeks to lose the same amount of weight that Joey loses in a single week. Elucidation: Sandy’s weight loss rate is slower than Joey’s weight loss rate. It takes Sandy 4 weeks to lose the same amount of weight that Joey loses in just one week. Segment 3: Joey loses 8 pounds in 4 weeks. Elucidation: Joey’s weight loss goal is to lose 8 pounds, and he achieves this goal in 4 weeks. Segment 4: How many weeks will it take Sandy to lose the same amount of weight? Elucidation: The question is asking for the number of weeks it will take Sandy to achieve the same weight loss goal as Joey. Solution: Since Joey loses 8 pounds in 4 weeks, his weight loss rate is 8/4 = 2 pounds per week. Since Sandy needs 4 weeks to lose the same amount of weight that Joey loses in a single week, her weight loss rate is 1/4 = 0.25 pounds per week. To lose 8 pounds like Joey, Sandy will need to lose at the same rate of 2 pounds per week. Therefore, it will take Sandy 8/2 = 4 weeks to lose the same amount of weight as Joey.
Question 3: Patty’s Plumbing charges \$40 to visit a house to make a repair, plus \$35 per hour, or part thereof, for labor, plus parts. One job took 2.25 hours and used \$60 in parts. How much did Patty charge?
CoT 1. Calculate the labor cost: The job took 2.25 hours, so Patty charged for 3 hours (since she charges for any part of an hour). The labor cost is \$35 per hour, so the labor cost for 3 hours is 3 x \$35 = \$105. 2. Calculate the total cost: The visit fee is \$40. The labor cost is \$105. The parts cost is \$60. The total cost is \$40 + \$105 + \$60 = \$205. Therefore, Patty charged \$205 for the job. (Correct)
@@
PEP + CoT: Segment 1: What is the base fee for Patty’s Plumbing to visit a house for a repair? Patty’s Plumbing charges \$40 to visit a house for a repair. Segment 2: How much does Patty’s Plumbing charge for labor? Patty’s Plumbing charges \$35 per hour, or part thereof, for labor. Segment 3: How much does Patty’s Plumbing charge for parts? The job used \$60 in parts. Segment 4: How long did the job take? The job took 2.25 hours. Solution: To calculate the total cost of the job, we need to add the base fee, labor cost, and parts cost. Base fee: \$40 Labor cost: \$35 per hour x 2.25 hours = \$78.75 Parts cost: \$60. Total cost: \$40 + \$78.75 + \$60 = \$178.75. Therefore, Patty charged \$178.75 for the job.